OpenStudy (anonymous):

Dan is planting 24 iries in his front yard.The flowers he bought were a combination of two varieties ,blue and white .The flowers are not blooming yet ,but Dan knows that the probability of having a blue flower is 75% 1)what is probability that at least 20 of white irises in Dan's garden? the key is 0.37

OpenStudy (lalaly):

Probability at least 20 white irises = P(X >= 20) = P(X = 20) + P(X = 21) + P(X = 22) + P(X = 23) + P(X = 24) = (24 C 20) (0.37)^20 (0.75)^4 + (24 C 21) (0.37)^21 (0.75)^3 + (24 C 22) (0.37)^22 (0.75)^2 + (24 C 23) (0.37)^23 (0.75)^1 + (24 C 24) (0.37)^24 (0.75)^0 =

OpenStudy (anonymous):

1-0.75=0.25 how you get 0.35 ?

OpenStudy (lalaly):

im sorry maybe i messd up wit calculations,,, u check them